# xcolorIncompatible color definition

### v2

@@ Abstract
Index Terms:
SAM, parameter efficient fine-tuning, LoRA, atrous attention, aortic vessel segmentation
I Introduction
@@
This paradigm shift has led to the development of advanced segmentation techniques, such as Segment Anything Model (SAM) [6], Swin-Unet [7], UNETR [8], SAMMed [9], and MedSAM [10], which leverage the power of ViT’s for highly accurate and computationally efficient segmentation tasks. The SAM enables users to generate segmentation masks through interactive prompts, such as clicks, bounding boxes, and text. Its exceptional zero-shot and few-shot capabilities have demonstrated strong effectiveness in natural image segmentation, garnering significant attention. However, despite the SAM’s success in natural image segmentation, recent studies have identified several limitations in its application to medical imaging [11][12]. These challenges stem from the inherent differences between natural and medical imaging data. Medical imaging datasets typically exhibit low contrast, ambiguous tissue boundaries, and small regions of interest. These limitations hinder SAM’s ability to generalize effectively without further fine-tuning [13].
@@
Recent studies [14, 15, 16] have sought to fine-tune SAM for medical image segmentation by incorporating domain-specific enhancements. However, fine-tuning these models demands substantial computational resources due to the large number of parameters in foundation models like SAM. Moreover, training large models on limited task-specific data frequently leads to overfitting and suboptimal performance. To address these challenges, parameter-efficient fine-tuning (PEFT) methods, such as Low-Rank Adaptation (LoRA) [17] have emerged as promising solutions. Several techniques have integrated LoRA into SAM to enhance computational efficiency while preserving performance, particularly in medical image segmentation [18][19].
@@
Despite these advancements, several fundamental intrinsic limitations of SAM persist. The SAM’s image encoder, based on plain ViTs, inherently lacks crucial vision-specific inductive biases needed to capture local patterns and fine details essential for dense predictions in medical imaging [20]. Additionally, SAM’s ViT-based architecture relies on global attention without integrating regional attention or sparse attention mechanisms, which are vital for focusing on relevant regions and reducing computational overhead [21]. Although regional attention captures spatial hierarchies at multiple scales, SAM’s reliance on global attention restricts its ability to focus on smaller, intricate regions in medical images. Moreover, the lack of sparse attention inhibits SAM from effectively capturing global context without incurring substantial computational costs. These limitations render SAM susceptible to errors, including the hallucination of small, disconnected components in segmentation [4][9], particularly when modeling structures such as vessels, tumors, or lesions. To enhance the performance of plain ViTs in dense prediction tasks, recent research has combined Transformer and convolutional features [22][23]. A study [24] integrates atrous attention with ViTs, enabling multi-scale feature extraction while preserving spatial resolution. Atrous attention unifies regional and sparse attention, enabling the model to focus on local details while simultaneously capturing the broader context.
@@
Figure 1: The architecture of the proposed VesselSAM framework. The model combines frozen pre-trained transformer blocks with learnable AtrousLoRA layers, enabling enhanced multi-scale feature extraction through Atrous Attention Module. Image embeddings generated by the frozen image encoder are fused with a bounding box prompt from the frozen prompt encoder and processed by the learnable mask decoder to produce the final segmentation. Blue represents frozen components, while orange denotes learnable parameters.
II Related Work
II-A ViT and SAM Based Medical Foundation Models
@@
II-B Parameter-Efficient Model Fine-Tuning
@@
The concept of Parameter-Efficient Fine-Tuning (PEFT) has emerged as an effective strategy to adapt large foundational models like SAM to specific downstream tasks with minimal additional parameter costs. One prominent PEFT approach, LoRA (Low-Rank Adaptation), has been successfully incorporated into SAM-based models. For instance,, SAMed[11] applied LoRA to SAM’s frozen image encoder, fine-tuning the LoRA layers, the prompt encoder, and the mask decoder together on medical datasets like Synapse multiorgan, demonstrating significant performance improvements. Similarly, SAMAdp[19] introduced a lightweight adapter module to enhance SAM’s segmentation performance in challenging tasks. By integrating task-specific prompts and adapters, SAMAdp improves segmentation accuracy while maintaining computational efficiency, demonstrating broad adaptability across diverse domains. Other works have pursued different approaches to optimize SAM for medical imaging applications. SAMMed[9] systematically evaluated SAM across 53 public medical imaging datasets, revealing that while SAM demonstrates strong zero-shot segmentation capabilities, its performance often degrades without fine-tuning, reinforcing the need for domain-specific adaptation.
II-C Atrous Convolution in ViTs
@@
Atrous Convolution (dilated convolution) has emerged as a powerful technique in Vision Transformers (ViTs) to enhance both local feature extraction and global contextual modeling, which are critical for segmentation tasks[25][26][27]. Atrous convolution expands the receptive field by introducing pixel “skipping”, enabling the model to capture multi-scale spatial dependencies without downsampling. This preserves fine-grained details while improving the ability to model broader spatial relationships. Initially introduced in DeepLab [20] for convolutional networks, Atrous Convolution has proven highly effective in extracting multi-scale features, which is crucial for handling segmentation tasks involving objects of varying sizes. In ViTs, where image features are typically processed as non-overlapping patches, integrating Atrous Convolutions enhances the model’s ability to learn hierarchical spatial dependencies. Specifically, Atrous Spatial Pyramid Pooling (ASPP) modules apply dilated convolutions at multiple rates, allowing the model to capture multi-scale contextual information [28], bridging the gap between local interactions and global dependencies. This approach is particularly beneficial in tasks requiring detailed segmentation, where capturing both local fine details and global context is necessary for accurate predictions. Recent advancements have shown that Atrous Convolutions are crucial for improving the performance of ViTs in segmentation tasks, particularly in domains such as medical imaging. In our model, we leverage the power of ASPP and Attention mechanisms to enhance the ViT encoder’s ability to capture both local priors and global context, effectively enabling the model to handle complex, high-resolution segmentation tasks with greater accuracy.
@@
Figure 2: LoRA and AtrousLoRA: A comparative overview with detailed insights into the Atrous Attention Module. Both LoRA and AtrousLoRA introduce a trainable encoder-decoder structure that operates in parallel with frozen pre-trained weights. (a) LoRA applies a low-rank constraint on the weight updates by factorizing them into smaller matrices (b) AtrousLoRA extends this approach by incorporating Atrous Attention Module into the bottleneck of LoRA, leveraging multi-scale dilated convolution operations for enhanced feature extraction. (c) The Atrous Attention Module features an Atrous Spatial Pyramid Pooling (ASPP) module with various dilation rates, global image-level pooling, and an attention mechanism that refines feature maps through element-wise multiplication with attention weights.
III Methodology
III-A Overview
@@ III-C VesselSAM
III-D LoRA and AtrousLoRA
LoRA [17] has emerged as a PEFT method, enabling task-specific adaptations of pre-trained models while significantly reducing computational and memory overhead. LoRA introduces low-rank trainable matrices to approximate weight updates, effectively bypassing the need to fine-tune the entire model (Fig. 2 (a)). Specifically, it adds two small matrices, $W_{b}$ and $W_{a}$, while keeping the original weights $W_{O}$ frozen during training. Given a pre-trained weight matrix $W_{O}\in\mathbb{R}^{C_{\text{out}}\times C_{\text{in}}}$, LoRA modifies the forward pass of the model as
@@ III-D LoRA and AtrousLoRA
While LoRA is highly efficient for adapting pre-trained models, it lacks the ability to explicitly capture multi-scale contextual information, which is critical for vision tasks such as image segmentation and dense prediction. To address this limitation, we introduced AtrousLoRA which incorporates atrous (dilated) convolutions into the LoRA framework (Fig. 2(b)). Atrous convolutions expand the receptive field of the model without increasing the number of parameters, enabling it to capture both local and global dependencies.
Mathematically, with AtrousLoRA Eq. 1 changes to:
@@
III-E Atrous Spatial Pyramid Pooling
@@
where $Z\in\mathbb{R}^{B\times C\times 1\times 1}$ represents the globally pooled feature map, summarizing the spatial information into a single vector per channel.
The outputs of the dilated convolutions $Y_{i}$ and the global average pooling $Z$ are then concatenated into a single feature map as expressed in Eq. 6.
@@
III-F Atrous Attention Module
The Atrous Attention Module (Fig. 2(c)) is introduced as a novel attention mechanism for Vision Transformers, designed to fuse regional and sparse attention effectively. This approach allows us to capture both global context and local detail with efficient computational complexity, while preserving the hierarchical information present in medical images. Inspired by atrous convolution [24], which expands the receptive field by skipping rows and columns in the input feature map without increasing the number of parameters. Atrous Attention enables VesselSAM to focus on relevant anatomical structures across multiple scales. The process is shown in Algorithm 1.
@@ III-F Atrous Attention Module
III-G Prompt Encoder And Mask Decoder
@@
III-H Loss Function and Evaluation Metrics
@@
TABLE I: Performance Comparison of our Proposed VesselSAM with other ViTBased and SAMbased Models on AVT dataset
@@
Note: Bold indicates the best results and underline denotes the second best results. ”Vanilla” refers to versions using pre-trained weights, while ”FT” indicates fine-tuned versions on AVT datasets.
TABLE II: Performance Comparison of our Proposed VesselSAM with other ViTBased and SAMbased Models on TBAD Dataset
@@
Note: Bold indicates the best results and underline denotes the second best results. ”Vanilla” refers to versions using pre-trained weights, while ”FT” indicates fine-tuned versions on TBAD dataset.
IV Experiments
IV-A Datasets
@@ IV-C Quantitative results
IV-C1 Quantitative Evaluation Results for AVT Dataset
@@
IV-C2 Quantitative Evaluation Results for TBAD Dataset
@@
Figure 3: Qualitative visual results on the AVT (Dongyang, Rider, Kits) and TBAD Datasets under Bounding Box Prompts. The first column shows the input images, followed by the ground truth (GT) in the second column. The subsequent columns show the results of various segmentation models: VesselSAM, SAM, MedSAM, SAMAdp, SAMed, and SAMMed. Each model’s output is overlaid with color-coded regions for true lumen (pink), false lumen (green), and the GT boundary line (yellow). The blue box represents the bounding box-prompt used for segmentation. The images have been zoomed in to enhance visibility.
IV-D Qualitative results
@@
Figure 4: Qualitative visual results on the AVT (Dongyang, Rider, Kits) and TBAD datasets using bounding box prompts. The first column shows the input images, followed by the ground truth (GT) in the second column. Subsequent columns show the results of the following configurations: ”VesselSAM* w/o AMM” refers to the VesselSAM model with the MedSAM backbone but without the Atrous Attention Module (AMM), while ”VesselSAM** w/o AMM” employs the SAM backbone without AMM. ”VesselSAM* with AMM” and ”VesselSAM** with AMM” incorporate AMM with the MedSAM and SAM backbones, respectively. The outputs are color-coded to highlight the true lumen (pink), the GT boundary line (yellow), and the bounding box prompt (blue). The images have been zoomed in to enhance visibility.
TABLE III: Ablation study of Atrous Attention Module
@@
Figure 5: The impact of Atrous Attention Module on validation loss over training epochs.
Figure 6: DSC VS Rank: Comparison of Dice Similarity Coefficients (DSC) for Aortic Vascular Tree (AVT) datasets (Dongyang, KiTs, Rider) and TBAD dataset across different LoRA ranks (2, 4, 16, 32, and 64), illustrating the performance stability and optimal rank selection for segmentation tasks.
IV-E Ablation Study
@@
IV-E1 Impact of the Backbone and Atrous Attention Module
@@
IV-E2 Impact of LoRA Rank
@@
IV-F Limitations and Future Work

### v3

@@ Abstract
keywords:
[…]
1 Introduction
@@
This paradigm shift has led to the development of advanced segmentation techniques, such as Segment Anything Model (SAM) [6], Swin-Unet [7], UNETR [8], SAMMed [9], and MedSAM [10], which leverage the power of ViT’s for highly accurate and computationally efficient segmentation tasks. The SAM enables users to generate segmentation masks through interactive prompts, such as clicks, bounding boxes, and text. Its exceptional zero-shot and few-shot capabilities have demonstrated strong effectiveness in natural image segmentation, garnering significant attention. However, despite the SAM’s success in natural image segmentation, recent studies have identified several limitations in its application to medical imaging [11, 12]. These challenges stem from the inherent differences between natural and medical imaging data. Medical imaging datasets typically exhibit low contrast, ambiguous tissue boundaries, and small regions of interest. These limitations hinder SAM’s ability to generalize effectively without further fine-tuning [13].
@@
Recent studies [14, 15, 16] have sought to fine-tune SAM for medical image segmentation by incorporating domain-specific enhancements. However, fine-tuning these models demands substantial computational resources due to the large number of parameters in foundation models like SAM. Moreover, training large models on limited task-specific data frequently leads to overfitting and suboptimal performance. To address these challenges, parameter-efficient fine-tuning (PEFT) methods, such as Low-Rank Adaptation (LoRA) [17] have emerged as promising solutions. Several techniques have integrated LoRA into SAM to enhance computational efficiency while preserving performance, particularly in medical image segmentation [18, 19].
@@
Despite these advancements, several fundamental intrinsic limitations of SAM persist. The SAM’s image encoder, based on plain ViTs, inherently lacks crucial vision-specific inductive biases needed to capture local patterns and fine details essential for dense predictions in medical imaging [20]. Additionally, SAM’s ViT-based architecture relies on global attention without integrating regional attention or sparse attention mechanisms, which are vital for focusing on relevant regions and reducing computational overhead [21]. Although regional attention captures spatial hierarchies at multiple scales, SAM’s reliance on global attention restricts its ability to focus on smaller, intricate regions in medical images. Moreover, the lack of sparse attention inhibits SAM from effectively capturing global context without incurring substantial computational costs. These limitations render SAM susceptible to errors, including the hallucination of small, disconnected components in segmentation [4, 9], particularly when modeling structures such as vessels, tumors, or lesions. To enhance the performance of plain ViTs in dense prediction tasks, recent research has combined Transformer and convolutional features [22, 23]. Recently, the work [24] integrates atrous attention with ViTs, enabling multi-scale feature extraction while preserving spatial resolution. Atrous attention unifies regional and sparse attention, enabling the model to focus on local details while simultaneously capturing the broader context.
@@
Fig. 1: The architecture of the proposed VesselSAM framework. The model combines frozen pre-trained transformer blocks with learnable AtrousLoRA layers, enabling enhanced multi-scale feature extraction through Atrous Attention Module. Image embeddings generated by the frozen image encoder are fused with a bounding box prompt from the frozen prompt encoder and processed by the learnable mask decoder to produce the final segmentation. Blue represents frozen components, while orange denotes learnable parameters.
2 Related Work
2.1 ViT and SAM Based Medical Foundation Models
@@
2.2 Parameter-Efficient Model Fine-Tuning
@@
The concept of Parameter-Efficient Fine-Tuning (PEFT) has emerged as an effective strategy to adapt large foundational models like SAM to specific downstream tasks with minimal additional parameter costs. One prominent PEFT approach, LoRA (Low-Rank Adaptation), has been successfully incorporated into SAM-based models. For instance, SAMed [11] applied LoRA to SAM’s frozen image encoder, fine-tuning the LoRA layers, the prompt encoder, and the mask decoder together on medical datasets like Synapse multiorgan, demonstrating significant performance improvements. Similarly, SAMAdp [19] introduced a lightweight adapter module to enhance SAM’s segmentation performance in challenging tasks. By integrating task-specific prompts and adapters, SAMAdp improves segmentation accuracy while maintaining computational efficiency, demonstrating broad adaptability across diverse domains. Other works have pursued different approaches to optimize SAM for medical imaging applications. SAMMed [9] systematically evaluated SAM across 53 public medical imaging datasets, revealing that while SAM demonstrates strong zero-shot segmentation capabilities, its performance often degrades without fine-tuning, reinforcing the need for domain-specific adaptation.
2.3 Atrous Convolution in ViTs
@@
Atrous Convolution (dilated convolution) has emerged as a powerful technique in Vision Transformers (ViTs) to enhance both local feature extraction and global contextual modeling, which are critical for segmentation tasks [25, 26, 27]. Atrous convolution expands the receptive field by introducing pixel “skipping”, enabling the model to capture multi-scale spatial dependencies without downsampling. This preserves fine-grained details while improving the ability to model broader spatial relationships. Initially introduced in DeepLab [20] for convolutional networks, Atrous Convolution has proven highly effective in extracting multi-scale features, which is crucial for handling segmentation tasks involving objects of varying sizes. In ViTs, where image features are typically processed as non-overlapping patches, integrating Atrous Convolutions enhances the model’s ability to learn hierarchical spatial dependencies. Specifically, Atrous Spatial Pyramid Pooling (ASPP) modules apply dilated convolutions at multiple rates, allowing the model to capture multi-scale contextual information [28], bridging the gap between local interactions and global dependencies. This approach is particularly beneficial in tasks requiring detailed segmentation, where capturing both local fine details and global context is necessary for accurate predictions. Recent advancements have shown that Atrous Convolutions are crucial for improving the performance of ViTs in segmentation tasks, particularly in domains such as medical imaging. In our model, we leverage the power of ASPP and Attention mechanisms to enhance the ViT encoder’s ability to capture both local priors and global context, effectively enabling the model to handle complex, high-resolution segmentation tasks with greater accuracy.
@@
Fig. 2: LoRA and AtrousLoRA: A comparative overview with detailed insights into the Atrous Attention Module. Both LoRA and AtrousLoRA introduce a trainable encoder-decoder structure that operates in parallel with frozen pre-trained weights. (a) LoRA applies a low-rank constraint on the weight updates by factorizing them into smaller matrices (b) AtrousLoRA extends this approach by incorporating Atrous Attention Module into the bottleneck of LoRA, leveraging multi-scale dilated convolution operations for enhanced feature extraction. (c) The Atrous Attention Module features an Atrous Spatial Pyramid Pooling (ASPP) module with various dilation rates, global image-level pooling, and an attention mechanism that refines feature maps through element-wise multiplication with attention weights.
3 Methodology
3.1 Overview
@@ 3.3 VesselSAM
3.4 LoRA and AtrousLoRA
LoRA [17] has emerged as a PEFT method, enabling task-specific adaptations of pre-trained models while significantly reducing computational and memory overhead. LoRA introduces low-rank trainable matrices to approximate weight updates, effectively bypassing the need to fine-tune the entire model Fig. 2 (a). Specifically, it adds two small matrices, $W_{b}$ and $W_{a}$, while keeping the original weights $W_{O}$ frozen during training. Given a pre-trained weight matrix $W_{O}\in\mathbb{R}^{C_{\text{out}}\times C_{\text{in}}}$, LoRA modifies the forward pass of the model as
@@ 3.4 LoRA and AtrousLoRA
While LoRA is highly efficient for adapting pre-trained models, it lacks the ability to explicitly capture multi-scale contextual information, which is critical for vision tasks such as image segmentation and dense prediction. To address this limitation, we introduced AtrousLoRA which incorporates atrous (dilated) convolutions into the LoRA framework Fig. 2(b). Atrous convolutions expand the receptive field of the model without increasing the number of parameters, enabling it to capture both local and global dependencies.
Mathematically, with AtrousLoRA Eq.(1) changes to:
@@
3.5 Atrous Spatial Pyramid Pooling
@@
where $Z\in\mathbb{R}^{B\times C\times 1\times 1}$ represents the globally pooled feature map, summarizing the spatial information into a single vector per channel. The outputs of the dilated convolutions $Y_{i}$ and the global average pooling $Z$ are then concatenated into a single feature map as expressed in Eq. 6.
@@
3.6 Atrous Attention Module
The Atrous Attention Module Fig. 2(c) is introduced as a novel attention mechanism for Vision Transformers, designed to fuse regional and sparse attention effectively. This approach allows us to capture both global context and local detail with efficient computational complexity, while preserving the hierarchical information present in medical images. Inspired by atrous convolution [24], which expands the receptive field by skipping rows and columns in the input feature map without increasing the number of parameters. Atrous Attention enables VesselSAM to focus on relevant anatomical structures across multiple scales. The process is shown in Algorithm 1.
@@ 3.6 Atrous Attention Module
3.7 Prompt Encoder And Mask Decoder
@@
3.8 Loss Function and Evaluation Metrics
@@
Table 1: Performance Comparison of our Proposed VesselSAM with other ViTBased and SAMbased Models on AVT dataset
@@
Note: Bold indicates the best results and underline denotes the second best results. "Vanilla" refers to versions using pre-trained weights, while "FT" indicates fine-tuned versions on AVT datasets.
Table 2: Performance Comparison of our Proposed VesselSAM with other ViTBased and SAMbased Models on TBAD Dataset
@@
Note: Bold indicates the best results and underline denotes the second best results. "Vanilla" refers to versions using pre-trained weights, while "FT" indicates fine-tuned versions on TBAD dataset.
4 Experiments
4.1 Datasets
@@ 4.3 Quantitative results
4.3.1 Quantitative Evaluation Results for AVT Dataset
@@
4.3.2 Quantitative Evaluation Results for TBAD Dataset
@@
Fig. 3: Qualitative visual results on the AVT (Dongyang, Rider, Kits) and TBAD Datasets under Bounding Box Prompts. The first column shows the input images, followed by the ground truth (GT) in the second column. The subsequent columns show the results of various segmentation models: VesselSAM, SAM, MedSAM, SAMAdp, SAMed, and SAMMed. Each model’s output is overlaid with color-coded regions for true lumen (pink), false lumen (green), and the GT boundary line (yellow). The blue box represents the bounding box-prompt used for segmentation. The images have been zoomed in to enhance visibility.
4.4 Qualitative results
@@
Fig. 4: Qualitative visual results on the AVT (Dongyang, Rider, Kits) and TBAD datasets using bounding box prompts. The first column shows the input images, followed by the ground truth (GT) in the second column. Subsequent columns show the results of the following configurations: "VesselSAM* w/o AMM" refers to the VesselSAM model with the MedSAM backbone but without the Atrous Attention Module (AMM), while "VesselSAM** w/o AMM" employs the SAM backbone without AMM. "VesselSAM* with AMM" and "VesselSAM** with AMM" incorporate AMM with the MedSAM and SAM backbones, respectively. The outputs are color-coded to highlight the true lumen (pink), the GT boundary line (yellow), and the bounding box prompt (blue). The images have been zoomed in to enhance visibility.
Table 3: Ablation study of Atrous Attention Module
@@
Fig. 5: The impact of Atrous Attention Module on validation loss over training epochs.
Fig. 6: DSC VS Rank: Comparison of Dice Similarity Coefficients (DSC) for Aortic Vascular Tree (AVT) datasets (Dongyang, KiTs, Rider) and TBAD dataset across different LoRA ranks (2, 4, 16, 32, and 64), illustrating the performance stability and optimal rank selection for segmentation tasks.
4.5 Ablation Study
@@
4.5.1 Impact of the Backbone and Atrous Attention Module
@@
4.5.2 Impact of LoRA Rank
@@
4.6 Limitations and Future Work
@@ 5 Conclusion
6 Declaration of Competing Interest
The authors declare that they have no conflicts of interest.
7 Acknowledgements
This work was supported by the Natural Science Foundation of Guangdong Province (No. 2023A1515010673), in part by the Shenzhen Science and Technology Innovation Bureau key project (No. JSGG20220831110400001, No. CJGJZD20230724093303007, KJZD20240903101259001), in part by Shenzhen Medical Research Fund (No. D2404001), in part by Shenzhen Engineering Laboratory for Diagnosis & Treatment Key Technologies of Interventional Surgical Robots (XMHT20220104009), and the Key Laboratory of Biomedical Imaging Science and System, CAS and the University Chinese Academy of Sciences and Alliance of International Science Organization (ANSO) through 2021A8017729012.